A gas sample occupies 7.73 L at 22 C. What is the pressure given that there are 1.16 mol of gas in the sample ? Answer in units of atm.
1
Expert's answer
2020-04-22T09:16:59-0400

From the Mendeleev-Clapeyron equation we can express the pressure:

"p•V = \nu•R•T" "\to" "p = (\nu•R•T)/V"

p = (1.16(mol)•8.31((N•m)/(mol•K))•295K)/7.73(m3) = 367.9 N/m2 = 0.03 atm

Answer: p = 0.03 atm